You will find how to measure CPI-U U.S. city average by expenditure category.    The two periods are September 2018 and September 2019. By what percent did this index change? Remember to show all work, starting with the formula and working through to the answer.
Answer #1

CPI U in US Expenditure Category for All Cities Average Unadjusted as per BLS website, is

September 2018= 252.439

September 2019= 256. 759

Thus percentage change = ((September's 2019- September 2018)/September 2018))×100

= ((256.759-252.439)/252.439)) × 100= 4.32×100/ 252.439= 1.7%.

Source : Economics news releases, Bureau of Labor Statistics.